OpenStudy (anonymous):

if a and b are two values of x lying between 0 and 360 which satisfy equation 6cosx+ 8sinx=9 then what will be the value of sin(a+b)

OpenStudy (anonymous):

\[\text{Sin}\left[\text{ArcCos}\left[\frac{1}{50} \left(27+4 \sqrt{19}\right)\right]+\text{ArcCos}\left[\frac{1}{50} \left(27-4 \sqrt{19}\right)\right]\right] \]\[\{x\to 0.476268, x\to 1.37832\} \]A plot is attached.